# ComSD: Balancing Behavioral Quality and Diversity in Unsupervised Skill Discovery

### v1

@@ Abstract
Learning diverse and qualified behaviors for utilization and adaptation without supervision is a key ability of intelligent creatures. Ideal unsupervised skill discovery methods are able to produce diverse and qualified skills in the absence of extrinsic reward, while the discovered skill set can efficiently adapt to downstream tasks in various ways. Maximizing the Mutual Information (MI) between skills and visited states can achieve ideal skill-conditioned behavior distillation in theory. However, it’s difficult for recent advanced methods to well balance behavioral quality (exploration) and diversity (exploitation) in practice, which may be attributed to the unreasonable MI estimation by their rigid intrinsic reward design. In this paper, we propose Contrastive multi-objectives Skill Discovery (ComSD) which tries to mitigate the quality-versus-diversity conflict of discovered behaviors through a more reasonable MI estimation and a dynamically weighted intrinsic reward. ComSD proposes to employ contrastive learning for a more reasonable estimation of skill-conditioned entropy in MI decomposition. In addition, a novel weighting mechanism is proposed to dynamically balance different entropy (in MI decomposition) estimations into a novel multi-objective intrinsic reward, to improve both skill diversity and quality. For challenging robot behavior discovery, ComSD can produce a qualified skill set consisting of diverse behaviors at different activity levels, which recent advanced methods cannot. On numerical evaluations, ComSD exhibits state-of-the-art adaptation performance, significantly outperforming recent advanced skill discovery methods across all skill combination tasks and most skill finetuning tasks. Codes will be released at https://github.com/liuxin0824/ComSD.
@@ 1 Introduction&Research Background
Reinforcement Learning (RL) has proven its effectiveness in learning useful task-specific skills in the presence of extrinsic rewards  (Mnih et al., 2015; Levine et al., 2016; Ding et al., 2021; Narvekar et al., 2017; Li et al., 2019). The success of unsupervised learning in computer vision (Chen et al., 2020; Caron et al., 2020) and natural language processing (Brown et al., 2020; Devlin et al., 2018) further benefits task-specific RL with complex input (Laskin et al., 2020; Kostrikov et al., 2020; Yarats et al., 2021a; Stooke et al., 2021) by improving representation learning. However, the agents trained with lots of efforts are always hard to generalize their knowledge to novel tasks due to the task-specific supervision of extrinsic reward (Stooke et al., 2021). In addition to the generalization, intelligent agents should also be able to explore environments and learn different useful behaviors without any extrinsic supervision, like human beings. For the reasons above, unsupervised skill discovery is proposed and becomes a novel research hotspot (Gregor et al., 2016; Eysenbach et al., 2018). As a branch of unsupervised RL  (Pathak et al., 2017; 2019; Burda et al., 2018), unsupervised skill discovery  (Park et al., 2021; 2023; Strouse et al., 2021) also designs task-agnostic rewards and achieves unsupervised pre-training with these rewards, which guarantees the task-agnostic downstream generalization. The main difference is that skill discovery requires extra input as conditions, named skill vectors. They aim to discover useful task-agnostic policies that are distinguishable by skill vectors. The discovered skills can be implemented for downstream tasks in various ways, including skill finetuning, skill combination, skill imitation, and so on.
Currently, one of the most popular and effective classes of skill discovery is based on Mutual Information (MI) maximization (Eysenbach et al., 2018; Sharma et al., 2019; Liu & Abbeel, 2021a; Yang et al., 2023). Most of them try to design intrinsic rewards to estimate and optimize the MI between skill vectors and visited states. With RL maximizing intrinsic reward expectation, correlations between skills and states are distilled, i.e., different useful skills are discovered. Recent MI-based methods exhibit considerable results in different fields, including map exploration (Campos et al., 2020), robotic manipulation (Plappert et al., 2018) and video games (Bellemare et al., 2013). However, when facing challenging robot locomotion where the space of state and action are continuous, complex, and non-linear, existing advanced methods can’t balance the behavioral quality (exploration) and diversity (exploitation) well. Specifically, they can either only learn different static postures of lazy exploration (Eysenbach et al., 2018; Liu & Abbeel, 2021a) or only produce highly dynamic behaviors that are homogeneous and indistinguishable (i.e., insufficient exploitation) (Laskin et al., 2022b), while an ideal robot behavior set should contain diverse behaviors at different activity levels, enabling efficient downstream adaptation in various ways. Some recent works have also noticed this issue (Yang et al., 2023), as we have. However, they only address it on 2D exploration problems but remain powerless for the difficult multi-joint robot behavior discovery.
Figure 1: ComSD discovers different useful unsupervised skills through a novel multi-objective intrinsic reward. Skill combination and skill finetuning together provide comprehensive skill evaluation, considering both behavioral quality and diversity. Our ComSD exhibits state-of-the-art adaptation performance on both kinds of downstream tasks, which recent advanced methods cannot.
We attribute the above quality-versus-diversity conflict to their unreasonable MI estimation and their rigid intrinsic reward design. In this paper, we propose ComSD, which tries to mitigate this conflict by a more reasonable MI estimator and a dynamic weighting algorithm for intrinsic reward design. Concretely, ComSD decomposes the MI between state transitions and skill vectors into the negative skill-conditioned entropy and state entropy, designing a novel intrinsic reward for MI estimation and optimization (with RL). For conditioned entropy, ComSD proposes to employ the contrastive learning result between skills and corresponding state transitions as a better estimation. For state entropy, ComSD follows recent works (Liu & Abbeel, 2021a; Yarats et al., 2021b; Laskin et al., 2022b), choosing a popular particle-based estimation (Liu & Abbeel, 2021b). Moreover, a novel Skill-based Multi-objective Weighting (SMW) mechanism is proposed to dynamically balance the above two entropy estimations into a novel multi-objective intrinsic reward, to encourage both exploration and exploitation. ComSD can produce a qualified skill set consisting of diverse behaviors at different activity levels for challenging multi-joint robots, which recent advanced methods cannot.
For comprehensive numerical evaluation, two different downstream adaptation tasks are employed: skill combination and skill finetuning. Skill finetuning is widely used recently (Laskin et al., 2022b; Yuan et al., 2022; Yang et al., 2023) on URLB (Laskin et al., 2021), where a target skill is chosen from the skill set and further finetuned with extrinsic reward. However, it’s one-sided to judge the whole skill set with only one skill, and the process of finetuning also introduces much uncertainty. In addition, behavioral diversity is also ignored in skill finetuning. To this end, we further employ another skill combination (Eysenbach et al., 2018) to evaluate both behavioral quality and diversity, where the learned skills are frozen and a meta-controller is trained to combine the learned skills for downstream task achievement. ComSD outperforms all baselines on skill combination significantly and is competitive with state-of-the-art methods on skill finetuning, as shown in Figure 1.
Our contributions can be summarized as follows: (i) We propose ComSD, a novel unsupervised skill discovery algorithm that discovers different useful behaviors by maximizing the MI objective between skill vectors and state transitions. (ii) ComSD proposes to employ the contrastive learning result between skills and states as a more reasonable conditioned entropy estimation in MI decomposition. (iii) A novel weighting algorithm, SMW, is presented to produce a novel multi-objective intrinsic reward that improves skill diversity and quality simultaneously. (iv) Comprehensive numerical evaluation and detailed analysis show that ComSD outperforms all recent advanced methods, exhibiting state-of-the-art downstream adaptation ability. It can produce a qualified skill set consisting of diverse behaviors for challenging multi-joint robots, which recent advanced methods cannot.
@@ 2.1 Problem Definition
Unsupervised skill discovery algorithms aim at discovering a useful and diverse set of agent behaviors in the absence of extrinsic reward, i.e., training a skill-conditioned policy in a task-agnostic environment. Concretely, a reward-free Markov Decision Process (MDP) (Bellman, 1957) is considered and defined as $\mathcal{M}^{free}=(\mathcal{S},\mathcal{A},\mathcal{P},\gamma,d_{0})$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $\mathcal{P}$ is the distribution of the next state given the current state and action, $\gamma$ is the discount factor, and $d_{0}$ is the distribution of the initial state. What skill discovery algorithms do is to define an intrinsic reward $r^{intr}$ and augment $\mathcal{M}^{free}$ to a intrinsic-reward MDP $\mathcal{M}^{intr}=(\mathcal{S},\mathcal{A},\mathcal{P},r^{intr},\gamma,d_{0})$. With $z\sim p(z)$ given by discovery algorithms, the skill-conditioned policy $\pi(a|s,z)$ can be obtained by RL over the $\mathcal{M}^{intr}$. As an example, we provide the pseudo-code of our ComSD in Appendix A.
To evaluate the adaption ability of discovered skills (i.e., the skill-conditioned agent $\pi(a|s,z)$), two adaptation evaluations: skill combination and skill finetuning are employed on each task-specific downstream task. A downstream task can be described as an extrinsic-reward MDP $\mathcal{M}^{extr}=(\mathcal{S},\mathcal{A},\mathcal{P},r^{extr},\gamma,d_{0})$, where $r^{extr}$ denotes the task-specific extrinsic reward. In skill combination, the learned $\pi(a|s,z)$ is frozen, and an meta-controller $\pi^{\prime}(z|s)$ is required to choose and combine skill vectors for $\pi(a|s,z)$ automatically on downstream tasks. Concretely, over the original task $\mathcal{M}^{extr}$, skill combination can be defined as $\mathcal{M}^{extr^{\prime}}=(\mathcal{S},\mathcal{Z},\mathcal{P^{\prime}},r^{extr^{\prime}},\gamma,d_{0})$, which regards different skill vectors $z\sim\mathcal{Z}$ as its actions and correspondingly changes its transition model and reward function. The meta-controller $\pi^{\prime}(z|s)$ is trained over $\mathcal{M}^{extr^{\prime}}$ by RL and then used for evaluation of the learned $\pi(a|s,z)$. In skill finetuning, a target skill vector $z_{i}$ is chosen. The corresponding policy $\pi(a|s,z_{i})$ is further finetuned on the $\mathcal{M}^{intr}$ for a few steps and serves as the evaluation of the discovered $\pi(a|s,z)$. The pseudo-codes of two adaptation tasks are provided in Appendix A.
@@ 2.2 Mutual Information Objective
As most skill discovery algorithms do, our ComSD also tries to maximize the Mutual Information (MI) objective between states $\tau(s)$ and skills $z$. In general, $\tau(s)$ can be (i) maintaining the original state $\tau(s)=s$, (ii) concatenating the neighboring state pairs $\tau(s)={\rm concat}(s^{t-1},s^{t})$, or (iii) using the whole trajectory $\tau(s)={\rm concat}(s^{1},...,s^{t})$. Following the recent advanced method (Laskin et al., 2022b), we define $\tau$ as state pairs $\tau(s)={\rm concat}(s^{t-1},s^{t})$ throughout this paper. The MI objective $I(\tau;z)$ can be decomposed into the form of Shannon entropy in two ways:
$I(\tau;z)=-H(z|\tau)+H(z)$ (1)
$I(\tau;z)=-H(\tau|z)+H(\tau)$ (2)
@@
where the $I(\cdot;\cdot)$ denotes the MI function and $H(\cdot)$ denotes the Shannon entropy throughout the paper. Most classical MI-based algorithms (Gregor et al., 2016; Achiam et al., 2018; Eysenbach et al., 2018; Lee et al., 2019)are based on the first decomposition Eq. 1. A uniform random policy can guarantee the maximum of the skill entropy $H(z)$ while the trainable discriminators estimating the conditioned entropy $H(z|\tau)$ are employed to calculate the intrinsic reward for unsupervised RL. However,  Laskin et al. (2021; 2022b) have found it hard for these methods to guarantee robot behavioral activity and exploration without external balancing signals. To this end, some recent works (Sharma et al., 2019; Liu & Abbeel, 2021a; Laskin et al., 2022b) focus on the second decomposition Eq. 2. They explicitly optimize the state entropy $H(\tau)$ by exploratory intrinsic rewards, achieving considerable performance on several tasks, including robotic manipulation, video games, and robot locomotion. Following these, we choose Eq. 2 as our optimization target.
3 Unsupervised Skill Discovery by ComSD
@@
Figure 2: Left: The multi-objective intrinsic reward design of ComSD. Right: In SMW, $\beta$ is linear related to $\rm flag\it(z)$ with different slopes in different region.
@@
In this section, we detail how ComSD designs a novel multi-objective intrinsic reward based on Eq. 2. In Section 3.1, we propose and describe how to employ contrastive learning to design an exploitation reward for the negative conditioned entropy estimation. In Section 3.2, we illustrate the way to estimate the state entropy by a particle-based exploration reward, which follows the recent advanced methods. In Section 3.3, we describe why and how to balance the above two sub-intrinsic rewards into a multi-objective intrinsic reward with proposed SMW. The overall intrinsic reward design of ComSD is visualized in Figure 2.
3.1 Conditioned Entropy Estimation via Contrastive Learning
@@
In decomposition Eq. 2, increasing the negative conditioned entropy $-H(\tau|z)$ intuitively encourages the agent to access familiar and visited state pairs according to the corresponding skills. This helps the agent exploit skill-specific information, i.e., increase the diversity of learned skills. Due to the unavailable conditioned distribution, we can’t directly maximize the $-H(\tau|z)$ but try to maximize its lower bound. First, $-H(\tau|z)$ can be decomposed as:
$-H(\tau|z)$ $=\sum_{\tau,z}p(\tau,z)\log p(\tau|z)=\sum_{\tau,z}p(\tau,z)\log\frac{p(\tau,z)}{p(z)}$ (3)
@@
$-H(\tau|z)$ $=\sum_{\tau,z}p(\tau,z)(\rm log\it p(\tau,z)-c)=\sum_{\tau,z}p(\tau,z)\rm log\it p(\tau,z)-\sum_{\tau,z}p(\tau,z)\cdot c$ (4)
$=\sum_{\tau,z}p(\tau,z)logp(\tau,z)-c\propto\sum_{\tau,z}p(\tau,z)\log p(\tau,z).$
Then we can obtain the variational lower bound:
$-H(\tau|z)$ $\propto\sum_{\tau,z}p(\tau,z)\log p(\tau,z)$ (5)
$=\sum_{\tau,z}p(\tau,z)\log p(\tau,z)+\sum_{\tau,z}p(\tau,z)\log q(\tau,z)-\sum_{\tau,z}p(\tau,z)\log q(\tau,z)$
$=\sum_{\tau,z}p(\tau,z)\log q(\tau,z)+D_{\mathrm{KL}}(p(\tau,z)||q(\tau,z))$
$=\mathbb{E}_{\tau,z}[\log q(\tau,z)]+D_{\mathrm{KL}}(p(\tau,z)||q(\tau,z))\geq\mathbb{E}_{\tau,z}[\log q(\tau,z)],$
where $q(\tau,z)$ is employed to estimate the unavailable $p(\tau,z)$. To optimize $-H(\tau|z)$, we need to minimize the KL divergence $D_{\mathrm{KL}}(p(\tau,z)||q(\tau,z))$ and maximize the expectation $\mathbb{E}_{\tau,z}[\log q(\tau,z)]$.
@@
where $f_{\theta_{1}}(\cdot)$ is the state encoder, $f_{\theta_{2}}(\cdot)$ is the skill encoder, T is the temperature, and $\exp(\cdot)$ makes $q$ non-negative like $p$. The above two encoders are updated by gradient descent after computing the NCE loss (Gutmann & Hyvärinen, 2010) in contrastive learning (Chen et al., 2020):
@@
For expectation maximization, we regard the optimization objective as a part of our unsupervised intrinsic reward, employing RL to optimize it:
$r^{intr}_{exploitation}\propto q(\tau_{i},z_{i}).$ (8)

### v2

# Balancing Both Diversity and Exploration in Unsupervised Discovery of Complex Behaviors

@@ Abstract
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
Unsupervised skill discovery seeks to dig out diverse and exploratory skills without extrinsic reward, with the discovered skills efficiently adapting to multiple downstream tasks in various ways. However, recent advanced methods struggle to well balance behavioral exploration and diversity, particularly when the agent dynamics are complex and potential skills are hard to discern (e.g., robot behavior discovery). In this paper, we propose Contrastive multi-objective Skill Discovery (ComSD) which discovers exploratory and diverse behaviors through a novel intrinsic incentive, named contrastive multi-objective reward. It contains a novel diversity reward based on contrastive learning to effectively drive agents to discern existing skills, and a particle-based exploration reward to access and learn new behaviors. Moreover, a novel dynamic weighting mechanism between the above two rewards is proposed for diversity-exploration balance, which further improves behavioral quality. Extensive experiments and analysis demonstrate that ComSD can generate diverse behaviors at different exploratory levels for complex multi-joint robots, enabling state-of-the-art performance across 32 challenging downstream adaptation tasks, which recent advanced methods cannot. Codes will be opened after publication.
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ 1 Introduction & Research Background
Under the supervision of extrinsic rewards, reinforcement learning (RL) enables agents to acquire task-specific skills  Mnih et al. (2015); Levine et al. (2016); Ding et al. (2021); Narvekar et al. (2017); Li et al. (2019). In contrast, real intelligent creatures, e.g., humans, are able to learn a number of useful task-agnostic behaviors without any supervision. To this end, unsupervised skill discovery is proposed to simulate this process for machines Gregor et al. (2016); Eysenbach et al. (2018); Park et al. (2021, 2023); Strouse et al. (2021); Shafiullah and Pinto (2022); Kim et al. (2023); Li et al. (2023); Zhao et al. (2022); Jiang et al. (2022). This approach aims to generate different useful skills through training a skill-conditioned policy with a designed intrinsic reward. The discovered skills should be as exploratory and diverse as possible, which enables efficient adaptation to multiple downstream tasks in various ways, including skill finetuning, skill combination, and so on  Eysenbach et al. (2018). Thus, both behavioral diversity and exploration are crucial for the discovered skills, and it’s significant to well balance them in unsupervised skill discovery.
Figure 1: Left: The complex agent dynamics and hard-to-distinguish potential behaviors significantly increase the difficulty of skill discovery. First, the difficulty of learning exploratory skills is increased. For example, learning to flip for a robot is much harder than reaching a far goal in map exploration due to the different agent dynamics. Second, the difficulty of discerning skills (i.e., promoting diversity) is increased. For example, reaching different goals can be easily represented by ending states, while dynamic flip and walk cannot be differentiated by ending states. In addition, behaviors at different exploratory levels are needed for robots (e.g., postures or different speeds) but are not considered in map. Right: ComSD’s pipeline. ComSD discovers exploratory and diverse behaviors through a novel contrastive multi-objective reward, showing state-of-the-art adaptation performance on both kinds of downstream tasks, which recent advanced methods cannot.
Currently, many methods have shown promising results in domains with simple agent dynamics and easily distinguishable potential skills, such as map exploration Campos et al. (2020), robotic arm reaching Plappert et al. (2018), and video games Bellemare et al. (2013). However, for more challenging domains where the agent dynamics are complex and the potential behaviors are hard to distinguish, like robot behavior discovery Laskin et al. (2021), the skill discovery difficulty is dramatically increased (see Figure 1 left), and recent advanced approaches cannot take both behavioral exploration and diversity into account. For example, an ideal robot skill set should contain a variety of behaviors at different activity levels, while recent methods can either only learn different static postures (low exploration)  Eysenbach et al. (2018); Liu and Abbeel (2021a) or only produce highly dynamic behaviors that are homogeneous (low diversity)  Laskin et al. (2022b). Some recent works have also noticed this issue Yang et al. (2023), but they only address it on maze exploration problems but remain powerless for more complex robots. This is because their intrinsic rewards don’t provide effective incentives to seek and discern far-reaching skills for complex agents, and they also ignore the exploration-diversity conflict (see Section 3.3) when potential behaviors are hard to discern.
In this paper, we propose Contrastive multi-objective Skill Discovery (ComSD) which seeks exploratory and diverse behaviors through a novel contrastive multi-objective reward. It designs strong incentives to respectively encourage diversity and exploration through contrastive learning, with a novel dynamic weight proposed for their balance. Concretely, this reward aims to maximize state-skill mutual information (MI), which is decomposed into skill-conditioned entropy (related to diversity) and state entropy (related to exploration). For skill-conditioned entropy optimization, we conduct state-skill contrastive learning, utilizing the contrastive results as a novel diversity reward. For state entropy maximization, we employ a particle-based estimation as the exploration reward. With contrastive learning, the above rewards achieve accurate entropy estimation, thus providing effective encouragement. In complex domains with hard-to-discern behaviors, we observe a serious mutually exclusive effect between the above two entropy optimization processes (see Section 3.3). Thus, a Skill-based Multi-objective Weighting (SMW) mechanism is proposed to dynamically weight the above rewards into the contrastive multi-objective reward for exploration-diversity balance. We evaluate the adaptation ability of discovered behaviors on 16 skill combination tasks and 16 skill finetuning tasks (task description in Section 2.1). ComSD enables state-of-the-art adaptation performance compared with five popular advanced methods (Figure 1 right). A detailed behavioral quality analysis is also provided, demonstrating that ComSD indeed generates qualified robot behaviors of both exploration and diversity, which recent advanced methods cannot.
Our contributions can be summarized as follows: (i) We propose ComSD to balance exploration and diversity in complex behavior discovery through a novel contrastive multi-objective reward containing a diversity reward and an exploration reward. (ii) We propose to utilize the state-skill contrastive results as a novel diversity reward for an effective diversity incentive. (iii) A novel dynamic weighting mechanism (SMW) is presented to weight the above two rewards for diversity-exploration balance, further improving behavioral quality. (iv) Extensive experiments and detailed analysis show that ComSD can discover diverse robot behaviors at different exploratory levels, enabling state-of-the-art downstream adaptation ability on 32 challenging tasks.
@@ 2.1 Problem Definition
Unsupervised skill discovery trains a skill-conditioned policy in a task-agnostic environment. Concretely, a reward-free Markov Decision Process (MDP) Bellman (1957) is considered and defined as $\mathcal{M}^{free}=(\mathcal{S},\mathcal{A},\mathcal{P},\gamma,d_{0})$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $\mathcal{P}$ is the distribution of the next state given the current state and action, $\gamma$ is the discount factor, and $d_{0}$ is the distribution of the initial state. What skill discovery algorithms do is to define an intrinsic reward $r^{intr}$ and augment $\mathcal{M}^{free}$ to a intrinsic-reward MDP $\mathcal{M}^{intr}=(\mathcal{S},\mathcal{A},\mathcal{P},r^{intr},\gamma,d_{0})$. With $z\sim p(z)$ given by discovery algorithms, the skill-conditioned policy $\pi(a|s,z)$ can be obtained by RL over the $\mathcal{M}^{intr}$. As an example, we provide the pseudo-code of our ComSD in Appendix A.
To evaluate the quality of discovered skills (i.e., the skill-conditioned agent $\pi(a|s,z)$), two adaptation evaluations: skill combination and skill finetuning are employed on each task-specific downstream task. A downstream task can be described as an extrinsic-reward MDP $\mathcal{M}^{extr}=(\mathcal{S},\mathcal{A},\mathcal{P},r^{extr},\gamma,d_{0})$, where $r^{extr}$ denotes the task-specific extrinsic reward. In skill combination, the learned $\pi(a|s,z)$ is frozen, and an meta-controller $\pi^{\prime}(z|s)$ is required to choose and combine skill vectors for $\pi(a|s,z)$ automatically on downstream tasks. Concretely, over the original task $\mathcal{M}^{extr}$, skill combination can be defined as $\mathcal{M}^{extr^{\prime}}=(\mathcal{S},\mathcal{Z},\mathcal{P^{\prime}},r^{extr^{\prime}},\gamma,d_{0})$, which regards different skill vectors $z\sim\mathcal{Z}$ as its actions and correspondingly changes its transition model and reward function. The meta-controller $\pi^{\prime}(z|s)$ is trained over $\mathcal{M}^{extr^{\prime}}$ by RL and then used for evaluation of the learned $\pi(a|s,z)$. In skill finetuning, a target skill vector $z_{i}$ is chosen. The corresponding policy $\pi(a|s,z_{i})$ is further finetuned on the $\mathcal{M}^{intr}$ for a few steps and serves as the evaluation of the discovered $\pi(a|s,z)$. The pseudo-codes of two adaptation tasks are provided in Appendix A.
@@ 2.2 Mutual Information Objective
As current state-of-the-art approaches do Laskin et al. (2022b); Liu and Abbeel (2021a); Eysenbach et al. (2018), our ComSD tries to maximize the Mutual Information (MI) objective between state function $\tau(s)$ and skill vector $z$ for behavior excavation. $\tau$ is defined as state transitions ($\tau(s)={\rm concat}(s^{t-1},s^{t})$) throughout this paper. The MI objective $I(\tau;z)$ can be decomposed into the form of Shannon entropy in two ways:
$I(\tau;z)=-H(z|\tau)+H(z),$ (1)
$I(\tau;z)=-H(\tau|z)+H(\tau),$ (2)
@@
where the $I(\cdot;\cdot)$ denotes the MI function and $H(\cdot)$ denotes the Shannon entropy throughout the paper. Most classical MI-based algorithms Gregor et al. (2016); Achiam et al. (2018); Eysenbach et al. (2018); Lee et al. (2019) are based on the first decomposition Eq. (1). A uniform random policy can guarantee the maximum of the skill entropy $H(z)$ while the trainable discriminators estimating the state-conditioned entropy $H(z|\tau)$ are employed as the intrinsic rewards for unsupervised RL. However,  Laskin et al. (2021, 2022b) have found it hard for these methods to guarantee robot behavioral exploration without any external balancing signals. To this end, some recent works Sharma et al. (2019); Liu and Abbeel (2021a); Laskin et al. (2022b) focus on the second decomposition Eq. (2). They explicitly optimize the state entropy $H(\tau)$ by designing exploration rewards, achieving considerable performance in different domains, including robotic manipulation, video games, and challenging robot locomotion. Following these, we choose Eq. (2) as our optimization target.
3 Behavior Discovery by ComSD
@@
Figure 2: Left: The contrastive multi-objective reward $r^{intr}_{ComSD}$ design. Right: In SMW, $\beta$ is linear related to $\rm flag\it(z)$ with different slopes in different region. The skill space is divided into different regions with different learning objectives.
@@
In this section, we show how ComSD seeks complex behaviors via a novel contrastive multi-objective reward $r^{intr}_{ComSD}$ based on Eq. (2). In Section 3.1, we propose a novel diversity reward $r^{contrast}_{diversity}$ through contrastive learning to estimate the conditioned entropy. In Section 3.2, we illustrate the way to estimate the state entropy by a particle-based exploration reward $r_{exploration}$. In Section 3.3, we detail why and how to balance the above two rewards into our contrastive multi-objective reward with the proposed SMW. The overall design of contrastive multi-objective reward is shown in Figure 2.
3.1 Skill-conditioned Entropy Estimation via Contrastive Learning
@@
In decomposition Eq. (2), increasing the negative conditioned entropy $-H(\tau|z)$ intuitively encourages the agent to access familiar and visited state pairs according to the corresponding skills. This helps the agent exploit skill-specific information, i.e., increase the diversity of learned skills. Due to the unavailable conditioned distribution, we can’t directly maximize the $-H(\tau|z)$ but try to maximize its lower bound. First, $-H(\tau|z)$ can be decomposed as:
$-H(\tau|z)$ $=\sum_{\tau,z}p(\tau,z)\log p(\tau|z)$ (3)
$=\sum_{\tau,z}p(\tau,z)\log\frac{p(\tau,z)}{p(z)}$
@@
$-H(\tau|z)$ $=\sum_{\tau,z}p(\tau,z)(\rm log\it p(\tau,z)-c)$ (4)
$=\sum_{\tau,z}p(\tau,z)\rm log\it p(\tau,z)-\sum_{\tau,z}p(\tau,z)\cdot c$
$=\sum_{\tau,z}p(\tau,z)logp(\tau,z)-c.$
We neglect $c$ and then obtain the variational lower bound:
$-H^{\prime}(\tau|z)$ $=\sum_{\tau,z}p(\tau,z)\log p(\tau,z)$ (5)
$=\sum_{\tau,z}p(\tau,z)\log q(\tau,z)+D_{\rm KL}(p(\tau,z)||q(\tau,z))$
$=\mathbb{E}_{\tau,z}[\log q(\tau,z)]+D_{\rm KL}(p(\tau,z)||q(\tau,z))$
$\geq\mathbb{E}_{\tau,z}[\log q(\tau,z)],$
where $q(\tau,z)$ is employed to estimate the unavailable $p(\tau,z)$. To optimize $-H(\tau|z)$, we need to minimize the KL divergence $D_{\rm KL}(p(\tau,z)||q(\tau,z))$ and maximize the expectation $\mathbb{E}_{\tau,z}[\log q(\tau,z)]$.
@@
where $f_{\theta_{1}}(\cdot)$ is the state encoder, $f_{\theta_{2}}(\cdot)$ is the skill encoder, T is the temperature, and $\exp(\cdot)$ makes $q$ non-negative like $p$. The above two encoders are updated by gradient descent after computing the NCE loss Gutmann and Hyvärinen (2010) in contrastive learning Chen et al. (2020):
@@
For expectation maximization, we regard the content (the contrastive result) as our novel contrastive diversity reward, employing RL to optimize it:
$r^{contrast}_{diversity}\propto q(\tau_{i},z_{i}).$ (8)